# MOOSE-Chem2: Exploring LLM Limits in Fine-Grained Scientific Hypothesis Discovery via Hierarchical Search

### v1

@@ Abstract
Large language models (LLMs) have shown promise in automating scientific hypothesis generation, yet existing approaches primarily yield coarse-grained hypotheses lacking critical methodological and experimental details. We introduce and formally define the novel task of fine-grained scientific hypothesis discovery, which entails generating detailed, experimentally actionable hypotheses from coarse initial research directions. We frame this as a combinatorial optimization problem and investigate the upper limits of LLMs’ capacity to solve it when maximally leveraged. Specifically, we explore four foundational questions: (1) how to best harness an LLM’s internal heuristics to formulate the fine-grained hypothesis it itself would judge as the most promising among all the possible hypotheses it might generate, based on its own internal scoring-thus defining a latent reward landscape over the hypothesis space; (2) whether such LLM-judged better hypotheses exhibit stronger alignment with ground-truth hypotheses; (3) whether shaping the reward landscape using an ensemble of diverse LLMs of similar capacity yields better outcomes than defining it with repeated instances of the strongest LLM among them; and (4) whether an ensemble of identical LLMs provides a more reliable reward landscape than a single LLM. To address these questions, we propose a hierarchical search method that incrementally proposes and integrates details into the hypothesis, progressing from general concepts to specific experimental configurations. We show that this hierarchical process smooths the reward landscape and enables more effective optimization. Empirical evaluations on a new benchmark of expert-annotated fine-grained hypotheses from recent chemistry literature show that our method consistently outperforms strong baselines.
@@ 1 Introduction
Large language models (LLMs) have increasingly been applied to assist scientific research (Luo et al., 2025; Cambria et al., 2023), with one of the most ambitious applications being the automated discovery of novel and valid scientific hypotheses. However, current methods produce hypotheses that are criticized for being overly coarse, lacking sufficient detail, offering simplistic suggestions, or omitting concrete implementation strategies (Wang et al., 2024; Hu et al., 2024; Si et al., 2024). Corresponding author.
@@
Formally, we define the task as generating a fine-grained hypothesis given a research background—comprising a research question and established methodologies—and a coarse-grained hypothesis direction. We show that fine-grained scientific hypothesis discovery is a combinatorial search problem, as it requires selecting and composing a coherent set of concrete details from a vast space of plausible options—making it particularly challenging in practice. The difficulty is compounded by the fact that scientific hypothesis discovery is an inherently out-of-domain (OOD) problem: the correctness of a hypothesis is fundamentally unknown at the time of formulation.
@@
A straightforward baseline is greedy search over the reward landscape. However, its non-convex and complex structure makes naive greedy strategies prone to poor local optima. To address this, we propose a hierarchical search framework that explicitly models how a finite-capacity reasoning agent—human or LLM—navigates the hypothesis space. Specifically, it first explores higher-level conceptual spaces and then incrementally refines into more specific detail spaces. This hierarchical approach smooths the reward landscape at each hierarchy level—especially at higher, more abstract levels—enabling convergence to superior local optima compared to greedy search and greedy search with self-consistency (Wang et al., 2023). The proposed framework inherently scales with the LLM’s available capacity, enabling systematic exploration of the limits of LLM-driven fine-grained hypothesis generation.
@@
Having investigated how to identify stronger local optima in Q1, we now turn to our second question (Q2): whether hypotheses judged better by LLMs exhibit stronger alignment with ground-truth hypotheses? To rigorously address Q2 while avoiding data contamination, we construct a benchmark of research backgrounds paired with expert-annotated fine-grained hypotheses from chemistry papers published after January 2024, ensuring these examples were unseen by our LLM (GPT-4o-mini, October 2023 cutoff). Using this benchmark, we indirectly evaluate Q2 by comparing the recall of hypotheses discovered by our hierarchical approach—which locates higher LLM-internal local optima—with hypotheses identified by baseline methods. Our results consistently show that hypotheses generated by our method achieve higher recall than those from baselines, providing empirical support for the reliability of the LLM’s internal reward signal in guiding fine-grained hypothesis discovery.
@@
Until now, the reward landscape guiding hypothesis search has been defined by a single LLM serving as the evaluator. We now turn to the third question (Q3): whether defining this landscape with an ensemble of diverse LLMs of similar capacity yields better outcomes than using multiple instances of the strongest LLM within that group. Our experiments show that ensembles composed of repeated instances of the strongest LLM consistently outperform equally sized ensembles of diverse models, suggesting that peak model quality is more important than architectural diversity in this setting.
@@
Finally, we consider a fourth question (Q4): whether an ensemble of identical LLMs provides a more effective reward landscape than a single instance of the same model. While Q3 compares ensembles of different models, Q4 isolates the effect of aggregation alone by controlling for model identity. We find that even identical LLMs, when sampled independently and aggregated via summarization, yield a reward signal that better captures novelty without sacrificing overall quality—highlighting a subtle but important dimension in optimizing hypothesis discovery.
@@
    We introduce and formalize fine-grained scientific hypothesis discovery as a combinatorial optimization problem, and release a post-2024 chemistry benchmark with expert-annotated fine-grained hypotheses, explicitly designed to prevent data contamination for current LLMs.
@@
    We systematically investigate this task through four foundational research questions: (Q1) how to best leverage an LLM’s internal heuristics for fine-grained hypothesis generation; (Q2) whether LLM-preferred hypotheses align more closely with ground-truth expert hypotheses; (Q3) whether ensembles of diverse LLMs provide a better reward landscape compared to repeated use of the strongest single model; and (Q4) whether ensembles of identical LLMs offer a better reward landscape than a single instance of the same model.
@@
    We propose a hierarchical search method over levels of conceptual abstraction, which smooths the reward landscape and reduces search complexity at each hierarchy level. Empirically, it consistently outperforms strong baselines in both LLM self-evaluation, expert evaluation, and recall against annotated ground-truth hypotheses.
@@ 2.1 Background and Task Motivation
Yang et al. (2024b) assume that many chemistry hypotheses can be constructed from a research background $b$—typically including the research question and/or background survey—and a set of inspirations $i_{1},\ldots,i_{k}$, representing concepts or findings from the literature. It can be formulated as:
@@
In practice, however, most hypotheses $h$ generated from Equation 1 tend to be coarse-grained: while they form cohesive associations between $b$ and the $i$, they often lack clear hypothesis specification and the detailed experimental configurations required for direct implementation in a laboratory setting. Additionally, many such hypotheses contain redundant or tangential elements—either due to the inclusion of unnecessary inspirations or from noise present in the literature that is unrelated to the core knowledge intended for hypothesis construction.
@@ 2.2 Problem Formulation: Fine-Grained Hypothesis Generation as Combinatorial Search
Here, $\{h_{c},d_{1},\ldots,d_{m}\}$ denotes the meaningful integration of edits $d_{1},\ldots,d_{m}$ into $h_{c}$, resulting in a coherent, fine-grained hypothesis. Each $d$ represents either: (1) the addition of a fine-grained detail to a concept $i$ in $h_{c}$, or (2) the deletion of a redundant concept from $h_{c}$. We define two sets of edit candidates: $D^{+}$, consisting of all possible fine-grained details that can be added to concepts in $h_{c}$, and $D^{-}$, consisting of all concepts in $h_{c}$ that may be removed. The full candidate set is defined as $D=D^{+}\cup D^{-}$.
@@ 2.3 Algorithmic Motivation for Hierarchical Heuristic Search (HHS)
Figure 1: Hierarchies designed for chemistry.
@@
A practical approach to this combinatorial complexity is to use heuristics that approximate solutions rather than exhaustively searching for exact ones. This aligns with how chemists refine hypotheses: given that $h_{c}$ often represents an unexplored direction, explicit candidate details $d$ are rarely retrievable from existing databases. Instead, chemists often rely on domain knowledge and intuition to heuristically identify and iteratively refine plausible details.
Analogously, we propose to leverage LLMs’ internal heuristics to guide the search for $d$ at each hierarchical level. As LLMs advance, their heuristics—emerging from pretraining over extensive scientific corpora—increasingly approximate, and in some cases surpass, those of human experts. By developing methods that maximally exploit these heuristics for fine-grained hypothesis generation, we aim to provide scientists with progressively greater support as LLMs’ capacities grow.
@@ 2.5 LLM-Based Implementation of Hierarchical Heuristic Search
We implement HHS as an LLM-driven agentic process that directly follows the hierarchical factorization formalized in Equation 6. As shown in Figure 2, at each hierarchy level $i$, $H_{i-1}$ represents the accumulated edits from all previous levels, corresponding to $D^{*(<i)}$. Within the current level, $h_{prev}$ denotes the partial hypothesis incorporating the edits selected up to step $j{-}1$, i.e., $d^{*(i)}_{<j}$. The candidate set $D^{(i)}$ is not explicitly enumerated but emerges implicitly from the LLM’s internal heuristics, conditioned on the background $b$, the hypothesis direction $h_{c}$, and the edits selected so far.
@@
At each iteration, the Add one $d$ at level $i$ module prompts an LLM to propose an edit $d$ to $h_{prev}$, producing a candidate $h_{cur}$, which is then refined once for validity, novelty, and specificity. The $h_{cur}>h_{prev}$ module evaluates whether the new hypothesis improves upon the previous one via LLM-based pairwise comparison, serving as an internal gradient signal.
@@
Within each hierarchy level, we adapt the design of an evolutionary unit (Yang et al., 2024b) to our task, where the search for the local optimum $h_{i}$ is independently repeated multiple times (set to three in our implementation), yielding distinct local optima $h_{i}^{1}$, $h_{i}^{2}$, and $h_{i}^{3}$. These candidates are then passed to a recombination module, which integrates their complementary strengths to interpolate a potentially superior local optimum $h_{i}$ within the subspace spanned by ${h_{i}^{1},h_{i}^{2},h_{i}^{3}}$.
@@ 2.6 Theoretical Analysis: Smoothing Effects of Hierarchical Heuristic Search
Figure 3: The smoothing effect of hierarchy on the hypothesis space performance landscape.
A key observation is that a hypothesis candidate’s performance at a higher hierarchy level can be viewed as an aggregated estimate—approximating an average or soft maximum—of its lower-level subspace. For instance, when evaluating a coarse-grained concept like “hierarchical 3D copper,” the LLM may implicitly account for diverse fine-grained structural variants, some highly relevant, others ineffective. We hypothesize that the LLM’s higher-level assessment aggregates these outcomes, weighting promising variants within the broader distribution to produce an overall estimate of the concept’s expected potential.
Building on this observation, the hierarchical abstraction smooths the reward landscape at higher levels by attenuating local irregularities in the fine-grained space, as the performance of a point at a higher level can be interpreted as an approximate aggregation or average of the performance across its corresponding lower-level subspace. This effect is illustrated in Figure 3 (a simplified schematic projection into a 1D space). Consequently, direct search over the flat, non-hierarchical space tends to be highly rugged and non-convex, often leading to premature convergence to suboptimal local optima. In contrast, introducing hierarchical structure progressively smooths the landscape, enabling more stable and efficient optimization, particularly at higher levels.
This smoothing effect can also be interpreted in the frequency domain as a form of low-pass filtering, where high-frequency components of the landscape are attenuated, resulting in a spectral cutoff in the spatial frequency domain, as illustrated in Figure 4.
Figure 4: Spectrum of the performance landscape over the hypothesis space: hierarchical design as a low-pass filter attenuating high-frequency irregularities.
@@ 3.1 Benchmark Construction, LLM Selection, and Baselines
To our knowledge, no existing benchmark provides annotated fine-grained scientific hypotheses—detailed enough for direct experimental execution. A critical consideration in constructing such a benchmark is minimizing data contamination by ensuring the LLM has not been exposed to the annotated data during pretraining. We extend the TOMATO-Chem dataset (Yang et al., 2024b), which contains 51 chemistry papers published and made available online after January 2024 in leading journals such as Nature and Science. Each entry is annotated with a research background $b$ and a coarse-grained hypothesis $h_{c}$. For this study, two PhD-level chemists further annotated these examples with fine-grained hypotheses $h_{f}$, providing ground-truth references for evaluation. To rigorously avoid data contamination, all experiments are conducted using GPT-4o-mini, whose pretraining data cutoff is October 2023.
@@
3.2 Q1: How to Best Harness an LLM’s Internal Heuristics to Formulate the Fine-Grained Hypothesis It Itself Would Judge as the Most Promising Among All Possible Hypotheses It Might Generate?
We frame this question as an optimization problem: Given only a coarse-grained hypothesis $h_{c}$ as the starting point, and relying entirely on a single LLM, how can we navigate the hypothesis space to approach the global optimum of the performance landscape, as defined by this same LLM’s internal heuristics, where each optimization step consists of adding an edit $d$ to $h_{c}$? In this setting, the LLM plays a dual role: it serves both as the proposal generator, proposing candidate edits $d$ to formulate new hypotheses within the hypothesis space, and as the gradient provider, judging whether the new hypothesis improves upon the current one via its own internal heuristics (e.g., pairwise comparison).
While it is inherently infeasible to determine whether a found local optimum represents the global optimum, we can empirically compare local optima obtained by different methods within the same performance pairwise evaluation, and therefore check which one is closer to a global optimum.
As detailed in § 2, the hierarchical design of HHS offers two key advantages over flat search strategies: (1) less search space to propose each $d$ (from $D^{(i)}$, instead of $D$), and (2) smoothing the performance landscape progressively in the hypothesis space. Among these, the smoothing effect is particularly critical, as it reduces the risk of early convergence to suboptimal local optima and facilitates progress toward higher peaks in the LLM’s internal performance landscape.

### v2

@@ Abstract
Large language models (LLMs) have shown promise in automating scientific hypothesis generation, yet existing approaches primarily yield coarse-grained hypotheses lacking critical methodological and experimental details. We introduce and formally define the new task of fine-grained scientific hypothesis discovery, which entails generating detailed, experimentally actionable hypotheses from coarse initial research directions. We frame this as a combinatorial optimization problem and investigate the upper limits of LLMs’ capacity to solve it when maximally leveraged. Specifically, we explore four foundational questions: (1) how to best harness an LLM’s internal heuristics to formulate the fine-grained hypothesis it itself would judge as the most promising among all the possible hypotheses it might generate, based on its own internal scoring-thus defining a latent reward landscape over the hypothesis space; (2) whether such LLM-judged better hypotheses exhibit stronger alignment with ground-truth hypotheses; (3) whether shaping the reward landscape using an ensemble of diverse LLMs of similar capacity yields better outcomes than defining it with repeated instances of the strongest LLM among them; and (4) whether an ensemble of identical LLMs provides a more reliable reward landscape than a single LLM. To address these questions, we propose a hierarchical search method that incrementally proposes and integrates details into the hypothesis, progressing from general concepts to specific experimental configurations. We show that this hierarchical process smooths the reward landscape and enables more effective optimization. Empirical evaluations on a new benchmark of expert-annotated fine-grained hypotheses from recent literature show that our method consistently outperforms strong baselines. [^1] All code and data can be found in https://github.com/ZonglinY/MOOSE-Chem2Contribution during internship at Shanghai Artificial Intelligence Laboratory. Corresponding author.
@@ 1 Introduction
Large language models (LLMs) have increasingly been applied to assist scientific research (Luo et al., 2025), with one of the most ambitious applications being the automated discovery of novel and valid scientific hypotheses. However, current methods produce hypotheses that are criticized for being overly coarse, lacking sufficient detail, offering simplistic suggestions, or omitting concrete implementation strategies (Wang et al., 2024; Hu et al., 2024; Si et al., 2025).
@@
Formally, we define the task as generating a fine-grained hypothesis given a research background—comprising a research question and a survey of established methodologies—and a coarse-grained hypothesis direction. We show that fine-grained scientific hypothesis discovery is a combinatorial search problem, as it requires selecting and composing a coherent set of concrete details from a vast space of plausible options—making it particularly challenging in practice. The difficulty is compounded by the fact that scientific hypothesis discovery is an inherently out-of-domain (OOD) problem: the correctness of a hypothesis is fundamentally unknown at the time of formulation.
@@
A straightforward baseline is greedy search over the reward landscape. However, its non-convex and complex structure makes naive greedy strategies prone to poor local optima. To address this, we propose a hierarchical search framework that explicitly models how a finite-capacity reasoning agent—human or LLM—navigates the hypothesis space. Specifically, it first explores higher-level conceptual spaces and then incrementally refines into more specific detail spaces. This hierarchical approach smooths the reward landscape at each hierarchy level—especially at higher, more abstract levels—enabling convergence to superior local optima compared to greedy search and greedy search with self-consistency (Wang et al., 2023). The proposed framework naturally scales with the capability of the underlying LLM’s heuristics, yielding better optima as those heuristics become stronger.
@@
Having investigated how to identify stronger local optima in Q1, we now turn to our second question (Q2): whether hypotheses judged better by LLMs exhibit stronger alignment with ground-truth hypotheses? To rigorously address Q2 while avoiding data contamination, we construct a benchmark of research backgrounds paired with expert-annotated fine-grained hypotheses from chemistry papers published after January 2024, ensuring these examples were unseen by our LLM (GPT-4o-mini, October 2023 cutoff). Using this benchmark, we indirectly evaluate Q2 by comparing the recall of hypotheses discovered by our hierarchical approach—which locates better LLM-internal local optima—with hypotheses identified by baseline methods. Our results consistently show that hypotheses generated by our method achieve higher recall than those from baselines, providing empirical support for the reliability of the LLM’s internal reward signal in guiding fine-grained hypothesis discovery.
@@
Until now, the reward landscape guiding hypothesis search has been defined by a single LLM serving as the evaluator. We now address our third question (Q3): whether defining this landscape with an ensemble of diverse LLMs of similar capacity yields better outcomes than using equally sized ensembles of the strongest LLM within that group. Our experiments show that ensembles of repeated instances of the strongest LLM consistently outperform equally sized ensembles of diverse models, suggesting that peak model quality is more critical than model diversity in this setting.
@@
Finally, we consider a fourth question (Q4): whether an ensemble of identical LLMs provides better reward landscape than a single instance of the same LLM. While Q3 compares ensembles of different models, Q4 isolates the effect of aggregation alone by controlling for model identity. We find that even identical LLMs, when sampled independently and aggregated via summarization, yield a reward signal that better captures novelty without sacrificing overall quality—highlighting a subtle but important dimension in optimizing hypothesis discovery.
Notably, while our experiments focus on chemistry, the task formulation, methodology, and analysis of Q1-Q4 are discipline-agnostic. The only domain-specific component is the manually designed hierarchy (used by the methodology), which can be instantiated for each new discipline encountered.
@@
    We introduce and formalize the fine-grained scientific hypothesis discovery task as a combinatorial optimization problem, and release a post-2024 benchmark with expert-annotated fine-grained hypotheses, specifically designed to prevent data contamination.
@@
    We explore the limits of LLMs for fine-grained scientific hypothesis discovery by framing it as an optimization problem over a reward landscape defined by LLM heuristics, with pairwise comparisons serving as the gradient signal. We also propose a hierarchical heuristic search framework that theoretically smooths the reward landscape, reduces search complexity, and identifies superior local optimum by interpolating among discovered optima. Empirically, this framework consistently outperforms strong baselines in locating better local optima.
@@
    We analyze this optimization formalization through 4 foundational research questions.
@@ 2.1 Background and Task Motivation
Yang et al. (2025) assume that many chemistry hypotheses can be constructed from a research background bb—typically including the research question and/or background survey—and a set of inspirations $i_{1},\ldots,i_{k}$, representing concepts or findings from the literature. It can be formulated as:
@@
In practice, however, most hypotheses $h$ generated from Equation 1 tend to be coarse-grained: while they form cohesive associations between $b$ and the $i$, they often lack clear hypothesis specification and the detailed experimental configurations required for direct implementation in a laboratory setting. Additionally, many such hypotheses contain redundant elements—either due to the inclusion of unnecessary inspirations or from noise present in the literature that is unrelated to the core knowledge intended for hypothesis construction.
@@ 2.2 Problem Formulation: Fine-Grained Hypothesis Generation as Combinatorial Search
Here, $\{h_{c},d_{1},\ldots,d_{m}\}$ denotes the meaningful integration of edits $d_{1},\ldots,d_{m}$ into $h_{c}$, resulting in a coherent, fine-grained hypothesis. Each edit $d$ corresponds to either (1) the addition of a fine-grained detail to an existing concept $i$ or the introduction of a new concept $i$ into $h_{c}$, or (2) the deletion of a redundant detail or concept from $h_{c}$. We define two edit candidate sets: $D^{+}$, containing all details and concepts that may be added to $h_{c}$; and $D^{-}$, containing all details and concepts within $h_{c}$ that may be removed. The overall edit space is then given by $D=D^{+}\cup D^{-}.$
@@ 2.3 Algorithmic Motivation for Hierarchical Heuristic Search (HHS)
Figure 1: Hierarchies designed for chemistry and material science by PhD-level domain expert.
@@
A practical approach to this combinatorial complexity is to use heuristics that approximate solutions rather than exhaustively searching for exact ones. This aligns with how chemists refine hypotheses: given that $h_{c}$ often represents an unexplored direction, while individual details may be retrievable from existing databases, the complete set of details is rarely available. Instead, chemists often rely on domain knowledge and intuition to heuristically identify and progressively integrate plausible details.
Analogously, we propose to leverage LLMs’ internal heuristics to guide the search for $d$ at each hierarchical level. As LLMs advance, their heuristics—emerging from pretraining over extensive scientific corpora—will increasingly approximate, and may surpass, those of human experts. The proposed framework naturally scales with the strength of these heuristics, yielding increasingly better optima for fine-grained hypothesis discovery as LLM capabilities continue to grow.
@@ 2.5 LLM-Based Implementation of Hierarchical Heuristic Search
We implement HHS as an LLM-driven agentic process that directly follows the hierarchical factorization formalized in Equation 6. As shown in Figure 2, at each hierarchy level $i$, $H_{i-1}$ represents the accumulated edits from all previous levels, corresponding to $D^{*(<i)}$. Within the current level, $h_{prev}$ denotes the partial hypothesis incorporating the edits selected up to step j1j{-}1, i.e., $d^{*(i)}_{<j}$. The candidate set $D^{(i)}$ is not explicitly enumerated but emerges implicitly from the LLM’s internal heuristics, conditioned on the background $b$, the hypothesis direction $h_{c}$, and the edits selected so far.
@@
At each iteration, the “Add one $d$ at level ii” module prompts an LLM to propose an edit $d$ to $h_{prev}$, producing a candidate $h_{cur}$, which is then refined once for validity, novelty, and specificity. The $h_{cur}>h_{prev}$” module evaluates whether the new hypothesis improves upon the previous one via LLM-based pairwise comparison, serving as an internal gradient signal for hypothesis optimization.
@@
Within each hierarchy level, we adapt the design of an evolutionary unit (Yang et al., 2025) to our task, where the search for the local optimum $h_{i}$ is independently repeated multiple times (set to three in our implementation), yielding several local optima $h_{i}^{1}$, $h_{i}^{2}$, and $h_{i}^{3}$. These candidates are then passed to a recombination module, which integrates their complementary strengths to interpolate a potentially superior local optimum $h_{i}$ within the subspace spanned by $h_{i}^{1},h_{i}^{2},h_{i}^{3}$}.
@@ 2.6 Theoretical Analysis: Smoothing Effects of Hierarchical Heuristic Search
Figure 3: The smoothing effect of hierarchy on the reward landscape of hypothesis space.
A key observation is that a hypothesis candidate’s performance at lower (more abstract) hierarchy level can be viewed as an aggregated estimate—approximating an average or soft maximum—of its higher-level subspace (more concrete). For instance, when evaluating a coarse-grained (more abstract) concept like “hierarchical 3D copper,” the LLM may implicitly account for its diverse fine-grained (more concrete) structural variants, some highly relevant, others ineffective. We hypothesize that the LLM’s assessment to the coarse-grained concept is an aggregation of its fine-grained outcomes, weighting promising variants within the broader distribution to produce an overall estimate of the coarse-grained concept’s expected potential.
Building on this observation, the hierarchical abstraction smooths the reward landscape at lower levels by attenuating local irregularities in the fine-grained space, as the performance of a point at a lower level can be interpreted as an approximate aggregation or average of the performance across its corresponding higher-level subspace. This effect is illustrated in Figure 3 (a simplified schematic projection into a 1D space). Consequently, direct search over the flat, non-hierarchical space tends to be highly rugged and non-convex, often leading to premature convergence to suboptimal local optima. In contrast, introducing hierarchical structure progressively smooths the landscape, enabling more stable and efficient optimization, particularly at lowest levels.
This smoothing effect can also be interpreted in the frequency domain as a form of low-pass filtering, where high-frequency components of the landscape are attenuated, resulting in a (roughly) spectral cutoff in the spatial frequency domain, as illustrated in Figure 4.
Figure 4: Hierarchical design as a low-pass filtering over the spectrum of the reward landscape.
@@ 3.1 Benchmark Construction, LLM Selection, and Baselines
To our knowledge, no existing benchmark provides annotated fine-grained scientific hypotheses—detailed enough for direct experimental execution. We extend the TOMATO-Chem dataset (Yang et al., 2025), which includes 51 chemistry papers published after January 2024 in leading journals such as Nature and Science. Each entry contains a research background $b$ and a coarse-grained hypothesis $h_{c}$. We further annotated it with fine-grained hypotheses $h_{f}$, serving as ground-truth references created by two PhD-level chemists. To prevent data contamination, all experiments are conducted using GPT-4o-mini, whose pretraining data cutoff is October 2023.
@@
3.2 Q1: How to Best Harness an LLM’s Internal Heuristics to Formulate the Fine-Grained Hypothesis It Would Judge Most Promising Among All It Might Generate?
We frame this question as an optimization problem: Given only a coarse-grained hypothesis $h_{c}$ as the starting point, and relying entirely on a single LLM, how can we navigate the hypothesis space to approach the global optimum of the reward landscape, as defined by this same LLM’s internal heuristics, where each optimization step consists of adding an edit $d$ to $h_{c}$? In this setting, the LLM plays a dual role: it serves both as the proposal generator, proposing candidate edits $d$ to formulate new hypotheses within the hypothesis space, and as the gradient provider, judging whether the new hypothesis improves upon the current one via its own internal heuristics (e.g., pairwise comparison).
While it is inherently infeasible to determine whether a found local optimum represents the global optimum, we can empirically compare local optima obtained by different methods with the gradient provider (pairwise comparison), and therefore check which one is a better local optimum.
As detailed in § 2, the hierarchical design of HHS offers two key advantages over flat search strategies: (1) less search space to propose each $d$ (from $D^{(i)}$, instead of $D$), and (2) smoothing the reward landscape progressively in the hypothesis space. Among these, the smoothing effect is particularly critical, as it reduces the risk of early convergence to suboptimal local optima and facilitates progress toward higher peaks in the LLM’s internal reward landscape.